OpenStudy (anonymous):

A total of 312 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (texaschic101):

a = adult tickets s = student tickets a + s = 312 -- (there were a total of 312 tickets sold) s = 3a -- (student tickets sold was 3 times adult tickets) these are the equations we will use, and since we are looking for adult tickets, lets change the first equation to : s = 312 - a now we sub 312 - a in for s in the 2nd equation. s = 3a 312 - a = 3a --- add a to both sides 312 = 3a + a 312 = 4a -- divide by 4 78 = a so there were 78 adult tickets sold